IF f(x) = -3X + 6, FIND F^-1(X)
Answer by venugopalramana(3286)

IF f(x) = -3X + 6=Y SAY, FIND F^-1(X)
-3X=Y-6
3X=6-Y
X=(6-Y)/3=2-Y/3
F^-1(X)=2- F(X)/3